# Quantum-Coherent Regime of Programmable Dipolar Spin Ice

### 30 Mar 2026 v1

@@ Results
    The data used to generate the animations and Fig. 3 are included in the archive along with the code. Requests for raw, shot‑by‑shot data should be directed to K.G. (krzysztof.giergiel@gmail.com).
  • Materials availability
    Not applicable
  • Code availability
    The codes for this study are available at doi.org/10.25919/vraf-2p08. The archive contains all code used to run the experiments via the D-Wave Ocean platform, as well as the scripts for data analysis and plotting, along with the necessary data.
  • Author contribution
Appendix A Additional plasma experiments
This section gathers additional results of the interacting monopole plasma expansion.
@@ Appendix A Additional plasma experiments
Supplementary Material:
Quantum-Coherent Regime of Programmable Dipolar Spin Ice
Emergent Gauge Field and Magnetic Charges in Artificial Spin Ice
Spin configuration and coarse-grained field
@@
Ice rule as a divergence constraint
@@
Magnetic charges from ice-rule violations
@@
Entropic interaction between monopoles
@@
where $F$ is the corresponding coarse-grained free-energy functional, and $T$ is the temperature. The coefficient $K$ is a dimensionless stiffness (entropic stiffness in the Coulomb-phase description) $v_{\mathrm{cell}}$ is the volume of a lattice cell [castelnovo_debye-huckel_2011]. Let us introduce a positive monopole charge +Q+Q at position $\mathbf{r}_{1}$ and a negative charge Q-Q at position $\mathbf{r}_{2}$. The divergence constraint becomes
@@
Dipolar Interaction Between Spins
@@
Implementation of long range interaction term $J_{3}$
We can use additional connecting qubit-spin between two spins which are not natively connected. This qubit will be driven with much higher X field ensuring its dynamics are fast. The dynamics of our $Z_{2}$ model are slow - hence the connecting qubit is experiencing slow - annealing dynamics. It’s contribution to energy follows the ground state configuration (for this qubit). The fast Hamiltionian looks as follows:
@@
$\Delta E=\sqrt{H^{2}+J^{2}}-|H|\approx\frac{J^{2}}{2H}\text{~ for small J}$ (18)
@@
Figure 7: Top Left: Schematic view of the layout positions - dots, and connections - curved lines, of the flux qubits on the DWAVE device. Coloured in dots are the qubits used to represent the ASI, rest is grayed out to indicate theu are unused. Yellow and black correspond to the horizontal/vertical dipoles in the spin ice state, as in Fig. 1 in the main text. The red dots, represent the qubits mediating the $J_{3}$ interactions. Blue and white color of connections represents the connections sign. Top Right: A unit cell associated with a single vertex superimposed on the structure. Center Left: Highlighted in green: representative $J_{\perp}$ connections. Center Right: Cyan - $J_{\parallel}$ connections. Bottom Left: Two horizontal and one vertical $J_{3}$ connections highlighted in red. Bottom Right: One horizontal and one vertical $J_{4}$ connections highlighted.
Testing the $t_{eff}=\Gamma\Delta t$.
@@
Impact of the energy scale of charge conservation enforcement?
We tested to check that the dynamics do not depend significantly on the choice of energy scale, which enforces the Gauss condition around the chosen value. We scanned over the strength of the local connections: $J_{\perp}^{\prime}=J_{\parallel}^{\prime}=C$ enforcing the charge conservation condition, see Fig. 9. We observe the transport is slightly inhibited by increasing the energy scale $C$ and when set to high we likely couple outside the qubit Hilbert space of the lowest energy states, to higher flux states. We selected the value 0.45 as the second lowest value point in our parameter scan, where the plotted fraction of charge conserved runs became stable. The animation attached to the article shows a recent sample of the configurations , when scaling the strength of the Gauss-Law terms. These were obtained from Advantage2 version 1.11 (corresponding to later date than the results shown in main text). The breaking of charge conservation is evident there, as at low values additional defects are being generated, and at the high values increased annihilation of the monopole pair is observed, as relaxation to the defect free spin ice is occuring.
@@
Calibration
@@
$\begin{vmatrix}Vertex&a_{hor}&b_{1}&b_{2}&b_{3}&b_{4}&a_{ver}\\ Counts&88915&82466&82911&82716&82362&76546\\ Exptected&92938&78281&78281&78281&78281&92938\end{vmatrix}$
[…]
where the expected row corresponds to the counts based on the numerical results for $20x20$ lattice with open boundary conditions presented in [Coraux2024], scaled to the number of (non-defective) vertices in our lattice (499). The numbers describe the number of vertices out of 499000 (based on the 1000 states of the 499 active vertices). The $a$, vertices correspond to the two directions of the type-1 vertices, and $b$ are four different type-2 vertices. The type-3 and -4 vertices are named according to their charges (|2Q||2Q|,$|4Q|$). When we run small Dirac string experiments with and without calibration, we observed the same scaling of the charge movement exponents, within our shot noise.
Obtained vertex statistics show we are close to an extensively degenerate six-vertex model square ASI. This is in contrast to $J_{\perp}\neq J_{\parallel}$ case of generalized F-model. Anisotropic experiments are usually the simpler to realize, as for example the natural dipolar interactions in nano-magnetic platforms prefer the line order  [Wang2006], but even bias in the opposite direction results in sub-extensive entropy [Perrin2016]. The $J_{\perp}=J_{\parallel}$ highly degenerate regime is more interesting as it maximizes the number of ground states giving rise to the entropic effects that are characteristic of ices.

### 19 May 2026 v2

@@ Acknowledgements.
The data used to generate the animations and Fig. 3 are included in the archive along with the code. Requests for raw, shot‑by‑shot data should be directed to K.G. (krzysztof.giergiel@gmail.com). The codes for this study are available at doi.org/10.25919/vraf-2p08. The archive contains all code used to run the experiments via the D-Wave Ocean platform, as well as the scripts for data analysis and plotting, along with the necessary data.
@@
Appendix A Emergent Gauge Field and Magnetic Charges in Artificial Spin Ice
A.1 Spin configuration and coarse-grained field
@@
A.2 Ice rule as a divergence constraint
@@
A.3 Magnetic charges from ice-rule violations
@@
Appendix B Entropic interaction between monopoles
@@
where $F$ is the corresponding coarse-grained free-energy functional, and $T$ is the temperature. The coefficient $K$ is a dimensionless stiffness (entropic stiffness in the Coulomb-phase description) $v_{\mathrm{cell}}$ is the volume of a lattice cell [5]. Let us introduce a positive monopole charge +Q+Q at position $\mathbf{r}_{1}$ and a negative charge Q-Q at position $\mathbf{r}_{2}$. The divergence constraint becomes
@@
Qualitative comparison with expectations from $\mathbb{Z}_{2}$ lattice gauge theories suggests that the observed screening behavior, correlation decay, and absence of extended confining strings are consistent with a deconfined regime. We emphasize, however, that our measurements do not directly probe standard diagnostics such as Wilson loops or topological sector structure, and therefore do not constitute a complete identification of a $\mathbb{Z}_{2}$ spin liquid.
The observed dynamics provide direct access to the microscopic motion of spin-ice defects in a quantum-coherent regime of an artificial platform. While we do not yet have a quantitative theory for the measured transport exponents, the anomalous dynamics—already present in the nominally non-interacting regime—can be attributed to constraints of the Hilbert space, which generate effective entropic interactions between monopoles (see Supplementary Material for a coarse-grained derivation).
Figure 7: Top Left: Schematic view of the layout positions - dots, and connections - curved lines, of the flux qubits on the DWAVE device. Coloured in dots are the qubits used to represent the ASI, rest is grayed out to indicate they are unused. Yellow and black correspond to the horizontal/vertical dipoles in the spin ice state, as in Fig. 1 in the main text. The red dots, represent the qubits mediating the $J_{3}$ interactions. Blue and white color of connections represents the connections sign. Top Right: A unit cell associated with a single vertex superimposed on the structure. Center Left: Highlighted in green: representative $J_{\perp}$ connections. Center Right: Cyan - $J_{\parallel}$ connections. Bottom Left: Two horizontal and one vertical $J_{3}$ connections highlighted in red. Bottom Right: One horizontal and one vertical $J_{4}$ connections highlighted.
Appendix C Dipolar Interaction Between Spins
@@
Appendix D Implementation of long range interaction term $J_{3}$
We can use additional connecting qubit-spin between two spins which are not natively connected. This qubit will be driven with much higher X field ensuring its dynamics are fast. The dynamics of our $Z_{2}$ model are slow - hence the connecting qubit is experiencing slow - annealing dynamics. It’s contribution to energy follows the ground state configuration (for this qubit). The fast Hamiltonian looks as follows:
@@ Appendix D Implementation of long range interaction term $J_{3}$
$\Delta E=\sqrt{H^{2}+J^{2}}-|H|\approx\frac{J^{2}}{2H}\text{\penalty 10000\ for small J}$ (18)
@@
Appendix E Testing the $t_{eff}=\Gamma\Delta t$.
@@
Appendix F Impact of the energy scale of charge conservation enforcement
We tested to check that the dynamics do not depend significantly on the choice of energy scale, which enforces the Gauss condition around the chosen value. We scanned over the strength of the local connections: $J_{\perp}^{\prime}=J_{\parallel}^{\prime}=C$ enforcing the charge conservation condition, see Fig. 9. We observe the transport is slightly inhibited by increasing the energy scale $C$ and when set to high we likely couple outside the qubit Hilbert space of the lowest energy states, to higher flux states. We selected the value 0.45 as the second lowest value point in our parameter scan, where the plotted fraction of charge conserved runs became stable. The animation attached to the article shows a recent sample of the configurations , when scaling the strength of the Gauss-Law terms. These were obtained from Advantage2 version 1.11 (corresponding to later date than the results shown in main text). The breaking of charge conservation is evident there, as at low values additional defects are being generated, and at the high values increased annihilation of the monopole pair is observed, as relaxation to the defect free spin ice is occurring.
Appendix G Calibration
@@
$\begin{vmatrix}Vertex&a_{hor}&b_{1}&b_{2}&b_{3}&b_{4}&a_{ver}\\ Counts&88915&82466&82911&82716&82362&76546\\ Expected&92938&78281&78281&78281&78281&92938\end{vmatrix}$
$\begin{vmatrix}-4Q&-2Q&-2Q&-2Q&-2Q&2Q&2Q&2Q&2Q&4Q\\ 0&429&395&354&384&399&399&359&365&0\\ 0&0&0&0&0&0&0&0&0&0\end{vmatrix}$
where the expected row corresponds to the counts based on the numerical results for $20x20$ lattice with open boundary conditions presented in [8], scaled to the number of (non-defective) vertices in our lattice (499). The numbers describe the number of vertices out of 499000 (based on the 1000 states of the 499 active vertices). The $a$, vertices correspond to the two directions of the type-1 vertices, and $b$ are four different type-2 vertices. The type-3 and -4 vertices are named according to their charges (|2Q||2Q|,$|4Q|$). We run small Dirac string experiments with and without this bias calibration and observed the same scaling of the charge movement exponents, within the shot noise. This suggests the main results should be robust even without the calibration step.
Obtained vertex statistics show that we are close to an extensively degenerate six-vertex model square ASI. This is in contrast to $J_{\perp}\neq J_{\parallel}$ case of the generalized F-model. Anisotropic experiments are usually the simpler to realize, as for example the natural dipolar interactions in nano-magnetic platforms prefer the line order [34], but even bias in the opposite direction results in sub-extensive entropy [30]. The highly degenerate regime $J_{\perp}=J_{\parallel}$ is more interesting as it maximizes the number of ground states giving rise to the entropic effects that are characteristic of ices.